In how many ways can a team of 20 hockey
players be accommodated in 10 two-person
hotel rooms? Assume that the order of
assigning the rooms does not matter.

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

In how many ways can a team of 20 hockey
players be accommodated in 10 two-person
hotel rooms? Assume that the order of
assigning the rooms does not matter.
-------------
Ans: 20C2 = (20*19)/(1*2) = 10*19 = 190 ways
---------------
Cheers,
Stan H.
---------

Answer by ikleyn(52777)

.
This mathematical conception is called "Combinations".

On Combinations, see the lessons
    - Introduction to Combinations
    - PROOF of the formula on the number of Combinations
    - Problems on Combinations
    - OVERVIEW of lessons on Permutations and Combinations
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.